Do you know how to use the foil method to solve y=2(x-3)(x-4)?
Answer by stanbon(75887)

use the foil method to solve y=2(x-3)(x-4)?
-----
Use FOIL on (x-3)(x-4)
F: x*x = x^2
O: x*-4 = -4x
I: x*-3 = -3x
L: -3*-4 = 12
-----
Combine to get: x^2 -7x + 12
---------
Your Problem:
y = 2(x^2-7x+12)
-----
y = 2x^2-14x+24
==========================
Cheers,
Stan H.
==================